Chapter 14 Homework Exercise 14.62 - Enhanced - with Feedback 34 of 48 Review I Constants I Periodic Table The rate constantThe rate constant of a reaction at 33 ∘C was measured to be 5.6×10−2 s−1. You may want to reference (Pages 606 - 612) Section 14.6 while completing this problem.
Answer #1

Giren, K = 5.6 X10 S-! A: 1-2 *10 -1 Temperature in kelvin Temperature in °c + 273.15 T = (33+ 273-15 lk T= 306.15k Universal